A rocket lifts off the surface of Earth with a constant acceleration of 20 m/s2. How fast will the rocket be going 1 min later? 


1
Expert's answer
2021-02-28T07:27:31-0500
"v=v_0+at,""v(t=60\ s)=0+20\ \dfrac{m}{s^2}\cdot60\ s=1200\ \dfrac{m}{s}."